28gms of Nitrogen molecule reacts with 6gm of hydrogen gas to form ammonia. Which one is the limiting reagent?
1
Expert's answer
2020-12-08T05:33:24-0500

The balanced reaction equation:

N2 + 3 H2 = 2 NH3

The quantity of nitrogen:

28 g / (28 g/mol) = 1 mol

The quantity of hydrogen:

6 g / (2 g/mol) = 3 mol

The ratio is 1:3. The same as according to the balanced equation.

Answer:

There is no limiting reagent.

The reaction mixture is stochiometric.